# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
where $E(u_{t})\coloneqq E(\omega_{u_{t}}^{n}/V)\coloneqq(\mathcal{I}-\mathcal{J})(u_{t}),$ $D(u_{t})\coloneqq D(\omega_{u_{t}}^{n}/V|\mu_{0})$ and $(\boldsymbol{D}/\boldsymbol{E})(\boldsymbol{u})\coloneqq\infty$ if $\boldsymbol{E}(\boldsymbol{u})=0.$ We recall that, $E(u_{t})$ and $D(u_{t})$ are convex with respect to $t$ (by the proof of [21, Lemma 9.1] and [27, Thm 4.10], respectively) and thus $\boldsymbol{D}$ and $\boldsymbol{E}$ are well-defined. The following proposition shows, in particular, that $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
@@
We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1].$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
@@
When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38, 9]. In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant). Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ under the constraint that $\mu$ has vanishing moment. In contrast to [37, 38, 9], in this case it does not seem possible to use symmetric decreasing rearrangements, due to a phenomenon of spontaneous symmetry breaking (see Section 4.1.1).
@@
(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $SO(3)$-invariant probability measure on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
@@
Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g_{S^{2}}}$ wrt the metric $g_{S^{2}}$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e. when $\omega_{0}=V\omega_{\text{FS}}$ for V=(22w),V=(2-2w),
(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu,$
where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e. $G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1.$
In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
@@
It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional.. Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (thus, importantely, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2. We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
@@
(4.7) $\mathcal{D}_{\gamma}(u)\leq\frac{1}{\gamma}F_{\gamma}\left(\frac{\omega_{u}}{V}\right).$
In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e. $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.39, has the desired properties. Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment. Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.37, has the desired properties. Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant. ∎
@@
is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [7, Cor 2] are obtained by replacing $\varphi$ with $\varphi$:
@@
Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10). In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.8 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
@@
[72] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.8 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5). But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{\gamma}$ that will be established in Theorem 5.1.
@@ 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.31. The conjecture was settled in [48] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.8 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.31 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87]. Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.9):
@@
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.31
@@
i.e. the inequality is saturated when $\varphi$ is constant. Equivalently,
@@
by applying the inequality 4.10 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.10 indeed holds if $u$ satisfies Tian’s constraint. ∎
@@
$F_{\gamma}(\mu)\geq F_{\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
@@
Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.10 gives
$F_{1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
Since $\mathcal{J}(u)=E(\omega_{u}/V)$ this equivalently means that $F_{1}(\omega_{u}/V)-E(\omega_{u}/V)/2\geq 0.$ In other words, $F_{3/2}(\omega_{u}/V)\geq 0.$ Using that $F_{\gamma}(\mu)$ is decreasing with respect to $\gamma$ (since $E\geq 0)$ this proves the inequality in the corollary. To prove the last statement first note that it follows from the previous proposition, combined with the scaling argument in the proof of Lemma 4.4, that
@@
if $\gamma\varphi$ satisfies Tian’s orthogonality condition and $\gamma\varphi\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{\varphi}=0$ (using that, by homogeneity, $\varphi$ satisfies Tian’s orthogonality condition, i.e. $0=\int\boldsymbol{x}\varphi\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{\varphi},$ using integration by parts). Furthermore, the condition that $\gamma\varphi\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{\varphi}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption. We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$

### v3

@@ 3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
where $E(u_{t})\coloneqq E(\omega_{u_{t}}^{n}/V)\coloneqq(\mathcal{I}-\mathcal{J})(u_{t}),$ $D(u_{t})\coloneqq D(\omega_{u_{t}}^{n}/V|\mu_{0})$ and $(\boldsymbol{D}/\boldsymbol{E})(\boldsymbol{u})\coloneqq\infty$ if $\boldsymbol{E}(\boldsymbol{u})=0.$ We recall that, $E(u_{t})$ and $D(u_{t})$ are convex with respect to $t$ (by the proof of [22, Lemma 9.1] and [29, Thm 4.10], respectively) and thus $\boldsymbol{D}$ and $\boldsymbol{E}$ are well-defined. The following proposition shows, in particular, that $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
@@
We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1[.$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
@@
When $\gamma=1$ the inequality holds without the moment constraint, i.e. it holds on all of $\mathcal{P}(X)$ (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant). If moreover, $w=0,$ the corresponding inequality on $\mathcal{P}(X)$ is then known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $S^{2}$ [43, 44, 10], which is the critical (logarithmic) case of the HLS inequality (as recalled in Section 5.2). Thus Theorem 3.13 yields, in particular, an improvement of the logarithmic HLS inequality on $S^{2}$ - allowing any $\gamma\in]0,2[$ under the constraint that $\mu$ has vanishing moment - that will be further sharpened in Section 5. In contrast to [43, 44, 10], in this case, it does not seem possible to use symmetric decreasing rearrangements. This is related to a phenomenon of spontaneous symmetry breaking, discussed in Section 4.1.1.
@@
(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $\mathrm{SO}(3)$-invariant probability measure $\mu_{0}$ on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
@@
Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g}$ wrt the metric $g$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e. when $\omega_{0}=V\mu_{0}$ for V=(22w)V=(2-2w)
(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu=\frac{V}{2}\left\|\mu-\mu_{0}\right\|_{H^{-1}}^{2}=\frac{V}{2}\left\|(-\Delta_{g})^{1/2}(\mu-\mu_{0})\right\|_{L^{2}(S^{2},dV_{g})}^{2}$
where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e. $G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1$ and $\left\|\mu-\nu\right\|_{H^{-1}}^{2}$ is the Hilbert norm on the Hilbert space $H^{-1}(\mathbb{P}^{1})$ of signed measures on $\mathbb{P}^{1}$ with zero mean that is dual to the Hilbert space $H^{1}(\mathbb{P}^{1}):=L^{2,1}(\mathbb{P}^{1})/\mathbb{R},$ endowed with the Hilbert norm $\left\|v\right\|_{H^{1}}^{2}:=\int_{\mathbb{P}^{1}}dv\wedge d^{c}v.$
In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [8, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
@@
It was introduced in [64] (when $\Delta=0)$ [^3] In the notation of [64], $\mathcal{D}_{\gamma}(u)=F_{-\gamma}(u),$ but here the notation $F_{-\gamma}$ is reserved for the free energy functional.. Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (thus, importantely, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2. We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
@@
(4.7) $\mathcal{D}_{\gamma}(u)\leq\frac{1}{\gamma}F_{-\gamma}\left(\frac{\omega_{u}}{V}\right).$
In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed,
(4.8) $F_{-\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$
using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e. $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.39, has the desired properties. Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment. Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.37, has the desired properties. Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant. ∎
@@
is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [8, Cor 2] are obtained by replacing $\varphi$ with $\varphi$:
@@
Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10). In this case Aubin [8, Cor 2] showed that the constrained Moser–Trudinger inequality 4.9 holds for any $\gamma<2.$ Moreover, in [71] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
@@
[80] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.9 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{-\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5). But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{-\gamma}$ that will be established in Theorem 5.1.
@@ 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.30. The conjecture was settled in [61] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.9 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.30 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [105]. Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.10):
@@
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.30
@@
Equivalently,
@@
by applying the inequality 4.11 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.11 indeed holds if $u$ satisfies Tian’s constraint. ∎
@@
$F_{-\gamma}(\mu)\geq F_{-\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
@@
Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.11 gives
$F_{-1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
Since $\mathcal{J}(u)=E(\omega_{u}/V)$ this equivalently means that $F_{-1}(\omega_{u}/V)-E(\omega_{u}/V)/2\geq 0.$ In other words, $F_{-3/2}(\omega_{u}/V)\geq 0.$ Using that $F_{-\gamma}(\mu)$ is decreasing with respect to $\gamma$ (since $E\geq 0)$ this proves the inequality in the corollary. To prove the last statement first note that it follows from the previous proposition, combined with the scaling argument in the proof of Lemma 4.4, that
@@
if $\gamma\varphi$ satisfies Tian’s orthogonality condition and $\gamma\varphi\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{\varphi}=0$ (using that, by homogeneity, $\varphi$ satisfies Tian’s orthogonality condition, i.e. $0=\int\boldsymbol{x}\varphi\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{\varphi},$ using integration by parts). Furthermore, the condition that $\gamma\varphi\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{\varphi}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption. We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{-\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$